Tech Staff Meeting Notes June 26, 2007

Current Top Priorities are: Station Maintenance. Mud-Island. Two new GPS stations. (Also wish to get INV populated.)


There are four levels of priority.
PrioritiesWish to try to get these done (or make progress on) before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about


This task list as of September 12, 2006. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.


The Fiscal Fish Fry will be friday, June 29, 2007. There are a total of 51 people signed up. 36 regular, 13 student, and 2 no fish.


Priorities

Station Maintenance, NMSZ
Several stations repaired. A bunch of single channel problems remain.
Lennox receive hardware and preamps replaced. Radios aligned, etc, etc. Signal strength is much better. Might consider doing the same at nmad and mkta.
Need a dual mux assembly or two. Have the AMUX64T's. Pat will build up the rest of the hardware.
New Madrid high temp alarm went on the other day. The ceiling is hot. Greg thinks he knows where to get Hypalon for the top of the roof. This would insulate and waterproof and avoid having to repaint. Should also go a long way toward keeping the inside cool.
MKTA needs an air conditioner in the doghouse.
Need a remote reset at lnxt. Current owner of the transmitter may not be installing a reset similar to what wkno had so we may need to install our own phone line. Steve suggested we get dsl while we're at it for a dbb backup but the cheap dutchman isn't ready to do that yet.

I-40 Bridge.
Need an 8 inch extension for surface vault to get it above grade.
New 200' hole is 100' so far. Bill is keeping keeping us posted.
Looks like scream2ew doesn't work with the new dm24. We just get binary hyroglyphics and Jim reports that the status streams look different too. Al will test with scream2ew from ew 7.1 to make sure that also doesn't work. Mitch will then contact Bruce.
After much griping from Mitch, it looks like the best solution is to telemeter to the weigh station. From there we get on the TDOT fiber network. We telemeter from the TDOT POP at Boswell St (aka perimeter mall) to Richardson. A pc will be installed at richardson (unless there is a viable switch solution to transfer from private TDOT to CERI subnet). From there we hop on the CERI microwave and suck the data into anssworm. Hopefully the project has sufficient funds for the extra telemetry link (about $4k) and pc (about $1.5k). Chris is arranging to have a fence installed (after any pad, gravel or whatever is installed). Cost is $1100 installed (must be real silver). We don't want to do it ourselves since it has to tie into the existing dilapidated fence.
The number 3 recorder keeps losing gps lock.
The recorder disks need to be reformatted. This can't be done remotely. The constant read/write to the fat file systems is going to make this reformatting something that needs to be done periodically. Requires physically going under the bridge to do it. Dont ya just love dos and obsolete hardware?
The wireless resets seem to be getting noisy making it more tricky than normal to reset.

GPS
Greg, Bob Smalley, James, Chris, Bill, and Mitch met on June 25, 2007 to discuss GPS issues. This was done outside the staff meeting set for June 26 because otherwise the entire 6/26 meeting would have been spent on GPS issues.

We're in the process of getting James out of the station maintenance business. GPS stations will be treated as any of the seismic network stations. Mac will be the lead for routine maintenance. Data handling and processing is still done by Bob et al.

Greg is in charge of getting anything done to the GPS stations to Steinerize them. This is necessary to make sure we can maintain them without burning alot of time on reapplying bandaids. Note that they are in pretty good shape already. The primary hassle is communications.

Greg has a couple options for improving communications.
Get permission from TWRA to put up a 50 or 60 foot tower at New Markham. From there we can probably hit the New Madrid node. It would be best to make it a 2.4G aironet shot so that its just an extension of the DBB. This would get us real-time continuous access to LCHS, NWCC, RLAP, and the new one at New Markham. the TWRA site is a concentrator for the four.
Can probably hit the lnxt node from STLE.
Not sure what to do about MAIR and CJTR.
Mitch noted that on the DBB, it is preferable to transfer less data more frequently that more data less often. This helps avoid longer burst periods that could slow down the real-time continuous seismic data. The GPS data bandwidth isn't large but if you burst 24 hours worth, it could cause a bunch of dropped seisnet packets.

James noted that the fans and filters need to be replaced at all sites.

Bob made the below notes in May. Greg will take the lead on monument install (which may be simply assigning tasks) after successful bid is in hand. He'll let Bob know when the comm etc, plan is finalized.

5/30/07

Install 2 new sites
	Select exact site
		LCHS - done
		New Markham - approx
	Get bids for monument
		Someone to supervise monument installation
		Someone to buy remaining supplies for monument installation
		Someone to weld monument after driller installed
	Build 2 boxes
	Spec solar requirements
	Lightning protection
	Install - monument then GPS and comms equipment

Upgrade/maint remaining sites
	Install SCIGN domes on all sites (excuse to visit each one - plan so can do all 
other upgrades, etc. needed)
	Communicaiton at LHCS (immediately brings NWCC online)
		How to get power and internet (wire) to base of tower
		Internet service*
		Lightning protection
	Upgrade at RLAP
		Find burried stuff
		Trench
		Solar (5 watts for 2 weeks)
		Install box on tower
		Install comms on tower (need LCHS up and running first,may or may not 
make it back without repeat through New Markham)
		Lightning protection
	Communications at STLE
		Select antenna site at School
		Run cables
		Install Freewave
		Power and antenna at STLE
		Internet service*
		Lightning protection
	Communications at CJTR
		Replace modem
		Investigate how to get on internet (Freewave to somewhere)
		Pay for DSL phone line?
	Communications at MAIR
		Find internet point of presence where we can piggyback

PTGV, MACC, MCTY, PIGT, CVMS, HCES - OK for now.

* confirm that contact at internet point of presense understands that we need to connect to 
the GPS receiver/our router from outside  has to be a public interent address.

ETN

Station Maintenance

John was on vacation. Jim and Chris provided an update.
May trip was a success. Want to schedule one per month for the next 3 months and finish things up.
CCRT is fixed.
Flat gap is done.
Van Hill is done. Holly says this is a very good station.
Greenback E/W sensor is replaced.
Van Hill gate got bull dozed. Need to replace it when they're done working.
Next trip will be CCNC install and repair at GFM and TRYN. Tentative date is July 9-13.
MGNC interference magically dissappeared.
Still might be a problem at SMNC.
TZTN needs a gate.

Network Completion

The map shows stations with yellow triangles being new ones.
stationman-daysstatus
MGNC8done
CCNC8new
WSSR8new
TVNC8new
VHTN8done
RBNC4done
TRYN4done
GFM4done
SMNC4done (see note above)
PLVA2abandon
BHT8move (eventually)

Routine Maintenance Kudzu, battery check, etc, etc.


Second Tier Priorities

ANSS.
C2SC was back and is down again. CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready.
CVTN is awaiting a new ip.
NAIT is the only non-real-time station. Wish to get it online. Might be able to hit Richardson from there.
NWS has new ups. They would like us to plug into that and remove our ups from beneath the rack. Bill will take care of it.

Lenox Agreement. We have new contact info. Mitch will see if a new MOU is desired.

FDL. There is a new FDL Policy. .
Bill reports that his laptop works well with the equipment but the others do not. Steve will show him how to get the others set up like his.
Steve has boot floppies and instructions for downloading data from the multi-channel seismograph without using the NT domain (which is defunct). The images and instructions are also available at /gaia/data/fdl/strataview.boot.floppy.


As Time Permits

3890 Power Chris is going to make a diagram for 3890. Good luck son.

Generators P-plant was over the other week and, to our surprise, ran the generator (the new one). Greg informed them that it runs automatically every Monday morning. Not really a good idea to run it under no-load anyway. Doesn't really solve the need for routine mainenance like oil and filter changes, etc. Mitch will try to find out what their plans are.

Inventory Jim has all the broadband sensors defined and is working on additional equipment (e.g. isis) and the connections. Mitch will install Bill's account in mysql and inv so he can start building stations. Mitch needs to sit down with Jim and figure out what we need to have ISTI help us with. Sid will be available after about mid-May to finish up anything we need his help to take care of. Apparently the downloaded docs from guralp for the cmg5td's have the old +-4g contants. Mitch has paper documentation with the correct +-2g constants.

Steiner Brain Dump Seminar Marathon week while John was in Memphis. Completed 12 sessions so far. Can fit 1 hr/dvd and each dvd is 4.5 GB. Have about 12.5 hours so far. Would be nice to get these online. How public should they be?


Other

SLU BB Stations. UTM is roughed in. Wurth needs to complete.
Poplar Bluff is done. The E/W channel in the sensor is bad so it needs to be replaced. Mitch will find out the scn and get them to add it to the ew export.

Thin-wire aironets are no longer available. This will be a problem for the DBB. It will need to be addressed hopefully later rather than sooner.

All hardware is in house to replace the node computers in the coming year. We'll upgrade to ew v7.1 at the same time. This means we'll be using scnl instead of just scn. Steve, Al, and Mitch are working on the details (headaches) of that. Glasworm is the converter and is operational with the new switches.
Note that the new MTTN node pc is the exception. We're not ready for it to be scnlized yet.

MPHB. Withers needs to determine current orientation so we can pull it and orient correctly. Helicorders look strange. Need to verify the data. Recentered, data are fine now.